OpenStudy (devonhoward15):

arrange these numbers is from least to greatest... will fan and medal for correct anwser -5/6, -4/9, -3/5, -4/7

jhonyy9 (jhonyy9):

divide every fractions and after you get these results you can easy rearrange them from least to greatest - ok. ?

OpenStudy (devonhoward15):

top from bottom?

OpenStudy (mathstudent55):

\(-\dfrac{5}{6} = -5 \div 6\)

OpenStudy (mathstudent55):

Use a calculator to change every fraction into a decimal. Then compare the numbers. Remember that all numbers are negative.

OpenStudy (devonhoward15):

i got -4/9, -4/7, -3/5, -5/6

OpenStudy (devonhoward15):

@mathstudent55

OpenStudy (mathstudent55):

-5/6 = -0.8333... -4/9 = -0.444... -3/5 = -0.6 -4/7 = -0.57142857...

OpenStudy (mathstudent55):

If your numbers were all positive, you'd be correct, but they are negative. With negative numbers, a number that looks larger is really smaller.

OpenStudy (devonhoward15):

oh ok

OpenStudy (mathstudent55):

For example, 100 is greater than 5, but -100 is less than -5.

OpenStudy (devonhoward15):

i see

OpenStudy (mathstudent55):